# Evaluating Decision Optimality of Autonomous Driving via Metamorphic Testing

### v1

@@ Abstract.
Autonomous Driving System (ADS) testing is crucial in ADS development, with the current primary focus being on safety. However, the evaluation of non-safety-critical performance, particularly the ADS’s ability to make optimal decisions and produce optimal paths for autonomous vehicles (AVs), is equally vital to ensure the intelligence and reduce risks of AVs. Currently, there is little work dedicated to assessing ADSs’ optimal decision-making performance due to the lack of corresponding oracles and the difficulty in generating scenarios with non-optimal decisions. In this paper, we focus on evaluating the decision-making quality of an ADS and propose the first method for detecting non-optimal decision scenarios (NoDSs), where the ADS does not compute optimal paths for AVs. Firstly, to deal with the oracle problem, we propose a novel metamorphic relation (MR) aimed at exposing violations of optimal decisions. The MR identifies the property that the ADS should retain optimal decisions when the optimal path remains unaffected by non-invasive changes. Subsequently, we develop a new framework, Decictor, designed to generate NoDSs efficiently. Decictor comprises three main components: Non-invasive Mutation, MR Check, and Feedback. The Non-invasive Mutation ensures that the original optimal path in the mutated scenarios is not affected, while the MR Check is responsible for determining whether non-optimal decisions are made. To enhance the effectiveness of identifying NoDSs, we design a feedback metric that combines both spatial and temporal aspects of the AV’s movement. We evaluate Decictor on Baidu Apollo, an open-source and production-grade ADS. The experimental results validate the effectiveness of Decictor in detecting non-optimal decisions of ADSs. It generates 46.0 NoDSs from 4 initial scenarios, while the best-performing baseline only detects 19.7 NoDSs. Our work provides valuable and original insights into evaluating the non-safety-critical performance of ADSs.
1. Introduction
@@
Autonomous Driving Systems (ADSs) have been a revolutionary technology with the potential to transform our transportation system into an intelligent one. ADSs aim to enable vehicles to operate without human intervention, relying on a combination of different sensors (e.g., camera, radar, lidar, and GPS) and artificial intelligence algorithms to perceive the environment, make decisions, and navigate safely. Even though the development of ADSs has seen significant progress over the past few decades, it is still a great challenge to guarantee that the ADSs can satisfy all performance requirements under different situations due to the existing vulnerabilities in ADSs (Garcia et al., 2020). Therefore, before their real-world deployment, ADSs should be sufficiently tested (Lou et al., 2022).
@@
Usually, the requirements of ADSs can be classified as safety-critical and non-safety-critical. Safety-critical requirements are those essential for ensuring safe operations and performance of the autonomous vehicle and the environment. For example, an ADS should guarantee that the ego vehicle (i.e., the autonomous vehicle controlled by the ADS) should arrive at its destination without causing collisions or violating traffic rules. Non-safety-critical requirements refer to aspects that do not directly impact safety performance but are important for positive user experiences and optimal vehicle performance, such as motion efficiency, passenger comfort, and energy consumption.
Currently, various testing technologies have been proposed to evaluate the safety-critical requirements of ADSs (referred to as “safety testing”), such as data-driven methods (Zhang and Cai, 2023; Deng et al., 2022b; Bashetty et al., 2020; Paardekooper et al., 2019) and guided searching methods (Hildebrandt et al., 2023; Cheng et al., 2023; Li et al., 2020; Haq et al., 2022; Zhong et al., 2023; Gambi et al., 2019b; Aleti et al., 2022). They aim to generate safety-critical scenarios, under which the ego vehicle will cause safety issues, such as collisions and traffic rule violations. Some ADS testing works (Luo et al., 2021; Huai et al., 2023) consider simple non-safety-critical requirements, such as comfort measured by acceleration. However, to the best of our knowledge, there is scant research on evaluating the decision optimality of ADSs (referred to as “optimal decision testing”). This aspect is an important non-safety-critical requirement and is crucial in generating optimal paths and guaranteeing motion efficiency for the ego vehicle in complex traffic scenarios. This type of testing aims to explore scenarios where safety risks are not present but remain crucial for assessing the overall intelligence and decision quality of the ADS.
Figure 1. Examples of non-optimal decision scenarios. Green lines are expected optimal paths and Red lines are non-optimal paths generated by the ADS.
Figure 1 showcases two examples of non-safety-critical violations where the ADS does not generate optimal paths. Note that these violations are not safety violations, as they still enable the ego vehicle to reach its destination safely. An intelligent ADS is expected to accurately appraise the traffic conditions and select the most efficient path to reach its destination. However, a less intelligent ADS may fall short in accurately interpreting traffic situations and make non-optimal decisions, resulting in non-optimal paths. It might unnecessarily hold back for an extended period despite a clear intersection or choose an inefficient path, leading to unnecessary delays or safety risks. Hence, assessing the quality of decision-making in ADSs is of significant importance. Non-optimal decisions can impact the overall comfort, efficiency, and energy consumption of autonomous vehicles and may even introduce potential safety risks.
In this paper, we present the first dedicated study on optimal decision testing for ADSs, focusing specifically on the detection of non-optimal decisions, i.e., those completing motion tasks safely with non-optimal paths. However, testing decision optimality and identifying corresponding violations pose unique challenges: (1) One challenge arises from the lack of explicit test oracles in optimal decision testing. Safety-critical violations and simple non-safety-critical violations (i.e., comfort) can be easily identified using clear and quantified criteria. For example, comfort can be assessed through abnormal acceleration. However, in optimal decision testing, no such oracles exist for evaluating the optimality of the paths made by the ADS. Determining whether the ADS has selected an optimal path or made the best decisions to reach the destination in a scenario becomes challenging. (2) Another challenge emerges due to the infinite and unpredictable nature of road conditions (Zhong et al., 2023; Haq et al., 2022). Road scenarios featuring variable road users and environmental factors create an infinite scenario space, complicating the generation of scenarios that violate the optimal decision requirement, referred to “Non-optimal Decision Scenarios” (NoDSs). Consequently, efficiently identifying NoDSs becomes a crucial challenge.
@@
To address these challenges, we introduce a novel metamorphic testing method called Decictor. This method effectively and efficiently detects violations of optimal decisions and the corresponding NoDSs. ❶ To tackle the first challenge, we propose a metamorphic relation (MR) that is designed to expose scenarios with non-optimal paths. The MR is defined as: the ADS should consistently make optimal decisions within a scenario, even after undergoing certain non-invasive changes that do not affect the original driving path. The violation of the MR indicates that the ADS is not making optimal decisions. ❷ To tackle the second challenge with regards to effectively detecting NoDSs, we propose a search-based testing method comprising three main components: Non-invasive Mutation, MR Check, and Feedback. Given an “Optimal Decision Scenario” (ODS) where the ADS produces an optimal path, Decictor employs non-invasive mutation to modify the behaviors of other participants, which does not affect the original optimal path taken by the ego vehicle. After the mutation, an abstraction-based behavior comparison is performed to check the validity of the MR. This step involves comparing the ego vehicle’s driving paths between the ODS and the mutated scenario. If the driving paths are different in the two scenarios, it suggests that the ADS has deviated from making optimal decisions. To provide feedback about the deviations that guide the generation of NoDSs, Decictor introduces a novel fitness function that considers two crucial metrics. First, it aims to maximize the differences between the driving paths of the ego vehicle in the two scenarios, directly accentuating the non-optimal behaviors. Second, it quantifies the similarity of the ego’s behavior, measured by a vector encompassing velocity, acceleration, and orientation. This metric is essential as the behavior of the ego vehicle directly influences the path it takes.
@@
We have evaluated Decictor on the Baidu Apollo with its built-in SimControl (Baidu, 2019). Since there is no existing work focusing on optimal decision testing, we compared Decictor with two random methods, i.e., Random, which generates scenarios randomly, and Random-$\delta$, which applies our non-invasive mutation and random selection to generate scenarios, and three state-of-the-art safety testing methods, i.e., AVFuzzer (Li et al., 2020), SAMOTA (Haq et al., 2022), and BehAVExplor (Cheng et al., 2023) on four meticulously validated ODSs. The evaluation results reveal the effectiveness and efficiency of Decictor in detecting NoDSs. Moreover, the comparison results emphasize the necessity of Decictor for optimal decision testing. Further experimental results demonstrate the utility of the mutation operation and feedback mechanisms meticulously designed within Decictor.
@@
  1. (1)
    We are the first to investigate the problem of non-optimal decisions made by ADSs. Specifically, we propose a metamorphic relation that effectively exposes non-safety-critical issues related to decision-making in ADSs.
  2. (2)
@@
    We develop a search-based algorithm that efficiently generates NoDSs, offering insights into non-safety-critical violations in ADSs.
  3. (3)
@@
    We conduct extensive experiments to evaluate the effectiveness and usefulness of Decictor on Baidu Apollo. As a result, a total of 46.0 NoDSs is discovered based on four initial ODSs on average, demonstrating the potential and value of Decictor in detecting non-optimal decisions made by ADSs.
2. Background and Notation
2.1. Autonomous Driving Systems
ADSs control the behavior of autonomous vehicles, i.e., the ego vehicles. Existing ADSs mainly contain two categories: End-to-End (E2E) systems (Zhang et al., 2021; Hu et al., 2023), and module-based ADSs (Kato et al., 2018; Baidu, 2019). E2E systems use united deep learning models to generate control decisions from sensor data directly. Recently, the rapid development of Deep Learning and Large Models have led to high-performance E2E systems in close-loop datasets, such as UniADS (Hu et al., 2023) and OpenPilot (comma.ai, 2022). However, these E2E systems still perform poorly on unseen testing data, such as easily colliding with obstacles. In contrast, module-based ADSs have better performance in various scenarios.
@@ 2.1. Autonomous Driving Systems
A typical module-based ADS, such as Baidu Apollo (Baidu, 2019) and Autoware (Kato et al., 2018), usually consists of localization, perception, prediction, planning, and control modules to generate decisions from rich sensor data. The localization module provides the location of the ego vehicle by fusing multiple input data from GPS, IMU, and LiDAR sensors. The perception module takes camera images, LiDAR point clouds, and Radar signals as inputs to detect the surrounding environment (e.g., traffic lights) and objects (e.g. other vehicles and pedestrians) by mainly using deep neural networks. The prediction module is responsible for tracking and predicting the trajectories of all surrounding objects detected by the perception module. Given the results of perception and prediction modules, the planning module then generates a local collision-free trajectory for the ego vehicle. Finally, the control module converts the planned trajectory to vehicle control commands (e.g., steering, throttle, and braking) and sends them to the chassis of the vehicle. In this paper, we choose to test module-based ADSs on a simulation platform, where the ADS connects to a simulator via a communication bridge, receives sensor data from the simulator, and sends the control commands to the vehicle in the simulator.
2.2. Scenario
@@
ADS testing necessitates a collection of scenarios as inputs. Each scenario is characterized by a specific environment (e.g., road, weather, and illumination) and the scenery and objects (static obstacles, Non-Player Character (NPC) vehicles, and pedestrians). In essence, a complex scenario can be generated by combining relevant attributes from the Operational Design Domains (ODDs) (Thorn et al., 2018). However, it is impractical to encompass all attributes with all possible values due to the vastness of the attribute space. Consequently, existing studies select different subsets of attributes for specific testing purposes. In this paper, we focus on evaluating the non-safety-critical performance of the motion generated by the ADS under test. As a result, we primarily employ traffic cones as static obstacles and NPC vehicles as dynamic objects to formulate our scenarios.
@@
Scenario. A scenario can be described as a tuple $s=\{\mathcal{A},\mathcal{P}\}$, where $\mathcal{A}$ is the motion task of the ADS under test, including the start position and the destination, $\mathcal{P}$ is a finite set of participants, including the set of static obstacles and dynamic NPC vehicles. Scenario observation is a sequence of scenes, and each scene represents the states of the ego vehicle and other participants at a timestamp. Formally, given a scenario $s=\{\mathcal{A},\mathcal{P}\}$, its observation is denoted as $O(s)=\{s_{0},s_{1},\ldots,s_{k}\}$, where $k$ is the length of the observation and $s_{i}$ is a scene at timestamp $i$. In detail, $s_{i}=\{y_{i}^{0},y_{i}^{1},\ldots,y_{i}^{|\mathcal{P}|}\}$ where $y_{i}^{j}=\{p_{i}^{j},\theta_{i}^{j},v_{i}^{j},a_{i}^{j}\}$ denotes the waypoint of a participant $j\in\mathcal{P}$ at timestamp $i$, including the center position $p_{i}^{j}$, the heading $\theta_{i}^{j}$, the velocity $v_{i}^{j}$ and the acceleration $a_{i}^{j}$. A driving path of the participant $j$ can be defined as $\tau^{j}(s)=\{p_{0}^{j},\ldots,p_{k}^{j}\}$. By default, we use $\tau(s)$ to represent the driving path of the ego vehicle in the scenario $s$. Unless otherwise specified, the driving path in the following context refers to the path of the ego vehicle in $s$, i.e., $\tau(s)$.
3. Overview
Figure 2. Overview of Decictor.
3.1. Problem Definition
This paper deviates from conventional evaluation of ADSs that typically focus on safety-related requirements. Instead, we aim to assess the optimization capability of the decision-making process in ADSs. Given a scenario $s$ and its associated potential optimal path for the ego vehicle, denoted as $\tau^{*}(s)$, our optimal decision testing endeavors to determine the ADS’s ability to consistently navigate along $\tau^{*}(s)$.
Unfortunately, it is hard to evaluate the ego vehicle’s driving path $\tau(s)$ is an optimal one due to the lack of testing oracles; that is, we do not know $s$’s optimal path of $\tau^{*}(s)$. Despite the possibility of manually assessing the driving path in each scenario, it becomes overwhelming, especially considering the multitude of scenarios involved in optimal decision testing. To address this, we employ metamorphic testing to detect the basic decision strategies. The metamorphic relation (MR) is defined as follows:
Definition 0 (Metamorphic Relation).
Given an ODS $s=\{\mathcal{A},\mathcal{P}\}$, we formalize the metamorphic relation as:
(1) $\forall P\subseteq\mathbb{P},\epsilon(\tau(s),\tau(\delta(s,P)))=True$
where $\mathbb{P}(\supseteq\mathcal{P})$ is the set of all possible participants in a scenario, the function $\tau$ retrieves the path that the ego vehicle takes in a scenario, and $\delta$ is a mutation function that modifies the behaviors of participants in $P$. It is crucial to note that the mutation function $\delta$ is designed in such a manner that even in the new scenario (i.e., $s^{\prime}=\delta(s,P)$), the ego vehicle can still take the driving path in the scenario $s$. The function $\epsilon$ is an equality function that assesses the equivalence of two paths.
The MR essentially posits that, within a specified new scenario (i.e., $s^{\prime}=\delta(s,P)$), the ADS should exhibit the ability to select the original optimal path. NoDSs can be identified by detecting violations of the MR. Note that the decision optimality of an initial scenario $\hat{s}$ can be manually confirmed while the mutated scenarios can be automatically checked by the MR.
Challenges. Implementing the MR presents several challenges, particularly concerning the functions $\delta$ and $\epsilon$. Not all mutated scenarios, i.e., $\delta(s,P)$, are suitable for the MR. One challenge lies in ensuring the ego vehicle can still traverse the original optimal path in the mutated scenario. For instance, if obstacles are introduced into the original optimal path, the ego vehicle may select a different path for safety considerations. While this new path differs from the original (i.e., $\tau(s)$), it may still be optimal within the context of the new scenario. The second challenge involves determining the path equivalence of two scenarios. The behavior of the ego vehicle may differ slightly between a NoDS and its corresponding ODS, such as minor differences between two paths. Even if they are not strictly identical, both paths could be optimal. Therefore, defining the function $\epsilon$ presents another challenge. Furthermore, once the MR is implemented, efficiently generating NoDSs that violate the MR remains challenging. To address these issues, this paper focuses on the development of the MR and the test generation algorithm.
3.2. Approach Overview

### v3

# Decictor: Towards Evaluating the Robustness of Decision-Making in Autonomous Driving Systems

@@ Abstract
Autonomous Driving System (ADS) testing is crucial in ADS development, with the current primary focus being on safety. However, the evaluation of non-safety-critical performance, particularly the ADS’s ability to make optimal decisions and produce optimal paths for autonomous vehicles (AVs), is also vital to ensure the intelligence and reduce risks of AVs. Currently, there is little work dedicated to assessing the robustness of ADSs’ path-planning decisions (PPDs), i.e., whether an ADS can maintain the optimal PPD after an insignificant change in the environment. The key challenges include the lack of clear oracles for assessing PPD optimality and the difficulty in searching for scenarios that lead to non-optimal PPDs. To fill this gap, in this paper, we focus on evaluating the robustness of ADSs’ PPDs and propose the first method, Decictor, for generating non-optimal decision scenarios (NoDSs), where the ADS does not plan optimal paths for AVs. Decictor comprises three main components: Non-invasive Mutation, Consistency Check, and Feedback. To overcome the oracle challenge, Non-invasive Mutation is devised to implement conservative modifications, ensuring the preservation of the original optimal path in the mutated scenarios. Subsequently, the Consistency Check is applied to determine the presence of non-optimal PPDs by comparing the driving paths in the original and mutated scenarios. To deal with the challenge of large environment space, we design Feedback metrics that integrate spatial and temporal dimensions of the AV’s movement. These metrics are crucial for effectively steering the generation of NoDSs. Therefore, Decictor can generate NoDSs by generating new scenarios and then identifying NoDSs in the new scenarios. We evaluate Decictor on Baidu Apollo, an open-source and production-grade ADS. The experimental results validate the effectiveness of Decictor in detecting non-optimal PPDs of ADSs. It generates 63.9 NoDSs in total, while the best-performing baseline only detects 35.4 NoDSs.
footnotetext: 1 Corresponding author
I Introduction
@@
Autonomous Driving Systems (ADSs) have been a revolutionary technology with the potential to transform our transportation system into an intelligent one. ADSs aim to enable vehicles to operate without human intervention, relying on a combination of different sensors (e.g., camera, radar, lidar, and GPS) and artificial intelligence algorithms to perceive the environment, make decisions, and navigate safely. Even though the development of ADSs has seen significant progress over the past few decades, it is still a great challenge to guarantee that the ADSs can satisfy all performance requirements under different situations due to the existing vulnerabilities in ADSs [1]. Therefore, before their real-world deployment, ADSs should be sufficiently tested in all aspects [2].
Figure 1: Illustration of non-optimal PPDs. Solid lines are expected optimal paths. Dotted lines are non-optimal paths planned by the ADS.
@@
Usually, the requirements of ADSs can be classified as safety-critical and non-safety-critical. Safety-critical requirements are those essential for ensuring safe operations and performance of the autonomous vehicle. For example, an ADS should guarantee that the ego vehicle (i.e., the autonomous vehicle controlled by the ADS) should arrive at its destination without causing collisions or violating traffic rules. Non-safety-critical requirements refer to aspects that do not directly impact safety but are important for user experiences and optimal vehicle performance, such as motion efficiency, passenger comfort, and energy consumption.
Currently, various testing technologies have been proposed to evaluate the safety-critical requirements of ADSs (referred to as “safety testing”), such as data-driven methods [3, 4, 5, 6] and search-based methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. They aim to generate safety-critical scenarios, under which the ego vehicle will cause safety issues, such as collisions and traffic rule violations. To the best of our knowledge, there is limited research on evaluating the robustness of the path-planning decisions (PPDs) in ADSs, i.e., their ability to plan optimal paths in dynamic environments. PPD optimality is an important non-safety-critical property for guaranteeing the motion efficiency of the ego vehicle in complex traffic scenarios, but it may not be achieved by ADSs. Fig. 1 presents two illustrative examples that highlight PPD robustness issues, where the ADS fails to determine the optimal paths after minor environmental changes, such as placing two small obstacles on the lane boundaries that would not affect the optimal path (i.e., solid lines). Note that, in these changes, the ego vehicle still maintains a safe distance [18] (i.e., more than 0.5 meters from the obstacle) to pass the scenarios if it follows the optimal paths. The video [19] shows an example that Tesla FSD cannot make robust PPDs at a fork in a real-world road.
An ADS with limited intelligence might inaccurately assess traffic scenarios, resulting in suboptimal choices. Such decisions could cause prolonged delays or select inefficient routes, leading to unnecessary hold-ups or even elevating safety risks. Therefore, evaluating the robustness of PPD of ADSs is critically vital, as it affects the comfort and efficiency of autonomous vehicles and could pose additional safety concerns. Note that, in this paper, PPD robustness evaluation aims to explore scenarios where safety issues are not present, instead focusing on assessing the optimal PPD capabilities of ADS within dynamic driving environments.
Technically, drawing on the foundational principles established by existing research in robustness evaluation research [20, 21], the core objective of robustness evaluation is to assess a model’s capacity to maintain correct prediction following some perturbations. For instance, in image classification, an image is slightly modified (e.g., through the addition of noise) to test whether the model can still accurately classify the perturbed image. Gradient-based methods [20, 21] are commonly employed to efficiently calculate these perturbations. However, evaluating the PPD robustness of ADS presents unique challenges: ❶ Contrary to safety testing, which relies on explicit, measurable criteria (like collision detection or timeout occurrences) to identify safety-critical violations, the assessment of PPD optimality in ADS under new scenarios lacks clear oracles. While traditional robustness evaluations might control the extent of perturbations (e.g., the level of noise) to ensure the perturbed inputs match the truth labels of original inputs, the complexity inherent in AV scenarios complicates the control of scenario perturbations. ❷ The unpredictable and infinitely variable nature of driving conditions, influenced by diverse road users and environmental factors, complicates the search of “Non-optimal Decision Scenarios” (NoDSs). The conventional gradient-based method is difficult to be applied in these scenarios due to discontinuities in the input space and the logical constraints that perturbations must obey, such as adherence to physical laws and ensuring that obstacles remain within realistic bounds. ❸ The evaluation of robustness typically requires a method to check prediction consistency between the original input and its perturbed counterpart. In classification tasks, the consistency check is straightforwardly verified by comparing their predicted labels. However, in ADS, decisions are represented by planned paths, which introduces complexity into the process of determining their consistency.
@@
To address these challenges, we propose a novel testing method, called Decictor, comprising three main components Non-invasive Mutation, Feedback and Consistency Check. The innovative approach enables the effective and efficient identification of non-optimal PPDs and their associated NoDSs. Specifically, to address the first challenge ❶, we design a non-invasive mutation that conservatively modifies the behaviors of other participants within the driving environment in a way that the modifications would not affect the originally optimal path in the given “Optimal Decision Scenario” (ODS). For the second challenge ❷, we propose a fitness function designed to guide the generation of NoDSs. The function scores are based on two key metrics: firstly, it seeks to maximize the discrepancy in the driving paths between the ego vehicle in both the mutated and original scenarios, directly accentuating the non-optimal behaviors. Secondly, it assesses the variances between the scenarios from a broader perspective of ego behavior, measured by a vector encompassing velocity, acceleration, and orientation. The second metric is crucial as the behavior of the ego vehicle may directly affect its driving path. For the third challenge ❸, given the possible slight variations between the new driving path (in the mutated scenario) and the initially optimal path (in ODS), the direct path-level comparison is not reliable. To resolve this, we introduce an abstraction-based behavior comparison method. This technique evaluates the driving patterns in both scenarios through abstraction, where significant differences in routes indicate deviations from optimal PPDs by the ADS.
@@
We have evaluated Decictor on the Baidu Apollo with its built-in SimControl [22]. We compared Decictor with nine baseline methods: two random methods, i.e., Random, which generates scenarios randomly, and Random-$\delta$, which applies our non-invasive mutation and random selection to generate scenarios, along with seven state-of-the-art ADS testing methods, i.e., AVFuzzer [8], SAMOTA [10], BehAVExplor [9], DriveFuzz [13], scenoRITA [14], DoppelTest [15] and DeepCollision [17]. The evaluation results reveal the effectiveness and efficiency of Decictor in detecting NoDSs, demonstrating the path-planning robustness issues within the ADS. For example, Decictor generates a total of 63.9 NoDSs while the best-performing baseline only detects 35.4 NoDSs. Further experimental results demonstrate the usefulness of our mutation operation and feedback mechanisms meticulously designed within Decictor.
@@
  1. 1.
    We are the first to investigate the problem of non-optimal PPDs made by ADSs. Our work provides valuable and original insights into evaluating the non-safety-critical performance of ADSs.
  2. 2.
@@
    We develop a search-based method that efficiently generates NoDSs and evaluates the PPD robustness of ADSs.
  3. 3.
@@
    We conduct extensive experiments to evaluate the effectiveness and usefulness of Decictor on Baidu Apollo. As a result, a total of 63.9 NoDSs is discovered on six initial ODSs on average, demonstrating the potential and value of Decictor in detecting non-optimal PPDs made by ADSs.
II Background and Notation
II-A Autonomous Driving Systems
@@
ADSs control the motion of autonomous vehicles. Existing ADSs mainly contain two categories: End-to-End (E2E) systems [23, 24, 25], and module-based ADSs [22, 26, 27]. E2E systems use united deep learning models to generate control decisions from sensor data directly. Recently, the rapid development of Deep Learning have led to high-performance E2E systems in close-loop datasets, such as OpenPilot [25]. However, these E2E systems still perform poorly on unseen data. In contrast, module-based ADSs have better performance in various scenarios. A typical module-based ADS, such as Baidu Apollo [22], usually consists of localization, perception, prediction, planning, and control modules to generate decisions from rich sensor data. The localization module provides the location of the ego vehicle by fusing multiple input data from GPS, IMU, and LiDAR sensors. The perception module takes camera images, LiDAR point clouds, and Radar signals as inputs to detect the surrounding environment (e.g., traffic lights) and objects (e.g. other vehicles) by mainly using deep neural networks. The prediction module is responsible for tracking and predicting the trajectories of all surrounding objects detected by the perception module. Given the results of perception and prediction modules, the planning module then generates a local collision-free trajectory for the ego vehicle. Finally, the control module converts the planned trajectory to vehicle control commands (e.g., steering, throttle, and braking) and sends them to the chassis of the vehicle. In this paper, we choose to test module-based ADSs on a simulation platform, where the ADS connects to a simulator via a communication bridge. Same to previous works [9, 15, 14], the ADS receives perfect perception data from the simulator, and sends the control commands to the ego vehicle in the simulator.
II-B Scenario
@@
ADS testing necessitates a collection of scenarios as inputs. Each scenario is characterized by a specific environment (e.g., road), the scenery and objects (e.g., static obstacles and Non-Player Character (NPC) vehicles). Complex scenarios are generated by combining relevant attributes from the Operational Design Domains (ODDs) [28]. Due to the vast attribute space, covering all attributes with all possible values is impractical. Existing studies [8, 9, 10, 14, 15, 17, 16, 13] address this by selecting specific subsets of attributes. In this paper, we evaluate the non-safety-critical performance of ADS motion using scenarios formulated with traffic cones as static obstacles and NPC vehicles as dynamic objects.
@@
Therefore, a scenario can be described as a tuple $s=\{\mathcal{A},\mathcal{P}\}$, where $\mathcal{A}$ is the motion task of the ADS under test, including the start position and the destination, $\mathcal{P}$ is a finite set of participants, including the set of static obstacles and dynamic NPC vehicles. Note that in our paper, participants in the scenario do not include the ego vehicle controlled by the ADS. Scenario observation is a sequence of scenes within the execution of the scenario, and each scene represents the states of the ego vehicle and other participants at a timestamp. Formally, given a scenario $s=\{\mathcal{A},\mathcal{P}\}$, its observation is denoted as $O(s)=\{o_{0},o_{1},\ldots,o_{k}\}$, where $k$ is the length of the observation and $o_{i}$ is a scene at timestamp $i$. In detail, $o_{i}=\{y_{i}^{0},y_{i}^{1},\ldots,y_{i}^{|\mathcal{P}|}\}$ where $y_{i}^{j}=\{p_{i}^{j},\theta_{i}^{j},v_{i}^{j},a_{i}^{j}\}$ denotes the waypoint of a participant $j\in\mathcal{P}$ at timestamp $i$, including the center position $p_{i}^{j}$, the heading $\theta_{i}^{j}$, the velocity $v_{i}^{j}$ and the acceleration $a_{i}^{j}$. A driving path of the participant $j$ can be defined as $\tau^{j}(s)=\{p_{0}^{j},\ldots,p_{k}^{j}\}$. By default, we use $\tau(s)$ to represent the driving path of the ego vehicle in the scenario $s$. Unless otherwise specified, the driving path in the following context refers to the path of the ego vehicle in $s$, i.e., $\tau(s)$.
III Overview
Figure 2: Overview of Decictor.
III-A Problem Definition
This paper aims to assess the optimization capability of the path-planning process in ADSs. Given an ODS $s$ and its associated optimal path for the ego vehicle $\tau(s)$, PPD optimality testing aims to assess the robustness of ADS’s PPD, i.e., its ability to consistently navigate along $\tau(s)$. Following the robustness definitions of neural networks [29, 30, 31, 32], we define the PPD robustness as follows:
Definition 1 (Path-planning Decision Robustness).
[…]
$\forall s^{\prime}.||s-s^{\prime}||_{\delta}=True\Longrightarrow\epsilon(\tau(s),\tau(s^{\prime}))=True$ (1)
where $||\cdot||_{\delta}$ is used to constrain scenario mutation, ensuring that the mutation $\delta$ modifies a subset of participants $P\subseteq\mathcal{P}$ in $s$ to produce $s^{\prime}$, without compromising the optimality of the original driving path, i.e., $||s-s^{\prime}||_{\delta}=True$. The function $\epsilon$ serves as a consistency checking mechanism, evaluating whether the two driving paths executed by the ADS in $s$ and $s^{\prime}$ are equivalent.
The definition of PPD robustness stipulates that, following minor perturbations (i.e., $s^{\prime}=\delta(s,P)$), the ADS should maintain the capacity to select the original optimal driving path (i.e., $\pi(s)$). The main goal of the PPD robustness testing is to identify scenarios where the ADS can accomplish its task without encountering safety issues, yet does not manage to plan and execute the optimal driving path. Note that the PPD optimality of the seed ODS $s$ can be manually confirmed while the PPD optimality of mutated scenarios $s^{\prime}$ can be automatically checked by evaluating their consistency.
III-B Approach Overview